# External Prompt Features Enhanced Parameter-efficient Fine-tuning for Salient Object Detection

### v1

@@ 4.1 Experiment Setup
Implemented with PyTorch, ExPert is trained on DUTS [21] dataset with a batch size of eight using two V100 16G GPUs. The encoder and the decoder are initialized with the public released pre-trained weights of SegFormer and the other parameters are initialized randomly. The learning rate is set to 2e-4 and the weight decay is set to zero.
@@ 4.2.1 Quantitative Comparison
We compare our model’s salient masks with four representative state-of-the art models on five salient object detection datasets: DUTS-TR, DUT-OMRON, ECSSD, HKU-IS and PASCAL-S. Two CNN SOTA models EGNet [27] and U2Net [17] are considered together with two transformer SOTA models SelfReformer [26] and EVP [16]. The metrics are calculated under the same condition using the prediction mask of different models [^3] We use the public codes of SOD_Evaluation_Metrics to compute the metrics.. The prediction masks are all provided by the official release [^4] EVP’s official mask is 352*352 which is not the original size, we resize the prediction map of EVP to the size of ground truth and then compute the metrics..
The results are shown in Table. 1. Our method achieves the best performance across all the SOTA models on all five datasets, which demonstrates the effectiveness of ExPert and the potential of transformer backbone on salient object detection. For MAE metric, ExPert surpasses the second best SelfReformer by 0.0058 (around 21% improvement) in ECSSD dataset and reaches 0.0198 in HKU-IS dataset. The superiority on all other three metrics demonstrates that ExPert has stronger competence to segment the salient objects in an image. Regarding the size of trained parameters, ExPert is more parameter efficient than all SOTA models except EVP which has a similar number of parameters as ExPert. Therefore, ExPert realizes a good trade-off between performance and size of trained parameters. Fig. 4 shows the F-measure curve and the precision-recall curve of ExPert and 4 SOTA models on five datasets. It is observable in these curves that our model consistently outperforms all other models.
Figure 4: The F-measure curve and the precision-recall (PR) curve of ExPert and four SOTA models on five datasets.
@@ 4.2.2 Qualitative Comparison
In Fig. 5, we show the qualitative comparison between our model and SOTA models to give readers an intuitive comprehension. Compared to SOTA models, ExPert’s masks are more accurate in details and can distinguish some ambiguous scenarios like reflection in water in the second row. The head of the girl is similar to the tree of background in the first row while the hair and the cloth similar to the dark background are confusing in the fifth row. Four SOTA models can’t differentiate these nuances but ExPert can handle the ambiguity. In the first row, ExPert’s mask is even more accurate than the ground truth which contains all the hair. Moreover, the semantic information injection from BLIP aids ExPert to recognize the relationship between objects in an image. For example in the fourth row, the baggage on the ground are obviously related to the man. ExPert segments them out while some other SOTA models neglect the baggage or focus wrongly on the street lamp. Another example is the third row, the children in the car should be regarded as a whole with the car. Semantic information also assists ExPert in handling complex scenarios, such as shadow interference or color similarity. For instance, discerning the body of the deer in the shadow.
Figure 5: The qualitative results of ExPert and four SOTA models. From left to right are the image, the ground truth, ExPert’s masks, EVP’s masks, SelfReformer’s masks, U2Net’s masks and EGNet’s masks. Better visual effect for zooming in.
@@ 4.3 Ablation study
Tuning methods: To verify the effectiveness of E-adapter on fine-tuning models for salient object detection, the performances of different fine-tuning methods are evaluated on ECSSD dataset. As shown in Table. 2, full fine-tuning achieves best performance but requires large size of trained parameters. Training from scratch and head tuning require large size of trained parameters and perform much worse than the other two tuning methods. As for adapter tuning, the performances is close to the fully tuned version but requires much fewer parameters.
Single-scale and multi-scale backbone: To compare the performance of single-scale backbone and multi-scale backbone, all fine-tuning methods are used to fine-tune MiT-B4 SegFormer and two of them are used to fine-tune ViT-B/16. The decoder of ViT is the same as [29], more details can be found in our supplement file. In Table. 2, training from scratch and adapter tuning of ViT are worse than the counterparts of SegFormer because multi-scale features can extract more detailed information in the images. Therefore, ExPert uses multi-scale pre-trained models as the encoder backbone for their finer details.
@@
Adapter level: In [4], the adapter module is side connected inside the transformer block while in ExPert we use the block-level adapter, which is side connected between transformer blocks. In Table. 2, the block-level adapter performs better than the FFT-level adapter which demonstrates that in segmentation task the block-level adapter are more advantageous. We suppose that the self-attention layers are also important in image recognition and should also be covered by adapters.
Prompt feature: To verify the effectiveness of E-injector and different prompt features, we denote the baseline model as the multi-scale encoder with E-adapter but without E-injector. In Table. 3, we compared E-injector with five prompt features to the baseline. Injecting DINO features as prompt features performs better than the baseline, which shows that object-aware features from DINO can guide the encoder to focus more on salient objects. The ViT features injection is better than the DINO features injection, suggesting that pre-trained models’ semantic information can further boost SOD performance. Additionally, although multi-scale backbone can extract multi-scale salient features, its global layers are usually shallower than single-scale. This weakness can be repaired by the injection of ViT’s features which go through more layers of the full-size scale.
Compared to the injection of ViT’s features, the injection of BLIP’s vision feature performs better. Since BLIP is trained with image-text pair, the semantic recognition of BLIP is stronger than ViT which is trained with image-label pair. This indicates that semantic information is critical for model to detect salient objects. The BLIP+ injection version surpasses BLIP Injection version, verifying that our cross-attention interaction between BLIP’s image features and caption embeddings highlight successfully the salient regions. The best performance comes from the combination of ViT’s feature and the BLIP+ feature, which capture both the detailed large scale features and the rich semantic information.
@@
Table 3: The results on ECSSD datasets of different prompt features for E-adapter. BLIP+ represents the interacted features after cross-attention of BLIP. Baseline only uses the E-adapter in 3.3 without the E-injector.
@@ 5 Conclusion
Looking ahead, further enhancement of ExPert may include exploration of additional prompt features. It is possible to inject other prompt features such as color or texture information into the backbone. Moreover, ExPert’s paradigm could also be applied to other segmentation tasks, such as semantic segmentation and panoptic segmentation. We leave these possible directions for future research to study.
@@ 6 More details of ExPert
The input image size of our model is 384*384, if the original size is not 384*384, we resize the input images to the certain size. The 384*384 resolution is used for ViT-B/16 [6] and MiT-B4 version of SegFormer [22] both training and inferring.
We also tried the single-scale pre-trained model Vit-B/16 as the backbone of ExPert. Since ViT model is trained for classifcation task, the decoder should change to a matched one. A simple decoder is adopted for ViT which is drawn from SETR [29]. For $N(=12)$ ViT layers, $N/4$ layers are selected with an interval of 4 as the feature set ${Z_{s}}=\{f_{2},f_{5},f_{8},f_{11}\}$ where $f_{i}$ is the vision feature of $i_{th}$ layer. Let $\mathcal{U}()$ as the upsampling, $Cat()$ as the concatenate operation applied to all $f_{i}$ in $Z_{s}$ and $\mathcal{S}()$ as the Sigmoid Function. The Mask $M$ is generated as Eq. 8 where $F_{c}$ is the concatenated feature. $Conv2()$ represents two convolution layers with an activation layer and $Conv()$ represents a single convolution layer.
@@ 7 More details of datasets and metrics
Training dataset: We use DUTS [21] as our training dataset following the previous works. DUTS is the largest salient object detection benchmark that contains training dataset and test dataset, including respectively 10,553 iamges and 5,019 images. The training images are all from the ImageNet DET training or val sets with sizes of slight difference. We use the test set of DUTS as our validation and evaluation dataset.
Evaluation dataset: We used ECSSD [23], DUT-OMRON [24], HKU-IS [11] and PASCAL-S [13] as the evaluation datasets. ECSSD dataset contains 1,000 real-world images across different domains with salient objects. DUT-OMRON is a challenging dataset with 5,168 high quality images with complex background and various kinds of objects. HKU-IS dataset has 4,447 well-annotated images characterised by multiple objects in diverse scenarios. PASCAL-S is a subset of PASCAL VOC segmentation validation dataset, containing 850 images. Unless specified, we used ECSSD for our default evaluation datasets for ablation experiments.
Evaluation metric: Four metrics are adpoted for our model evaluation. The first metric is the mean absolute error (MAE) as shown in Eq. 9, which computes the average pixel difference between predicted maps and ground truth labels. The smaller value of MAE signifies better performance.
@@
The second metric is F-measure $F_{\beta}$ [1] which is obtained leveraging the precision and the recall as Eq. 10. The precision and recall are computed using different thresholds from 0 to 255 between the predicted map and the ground truth. The $\beta^{2}$ is set to 0.3 followed the former works. We use the maximum $F_{\beta}$ for performance of each method. The larger value of $F_{\beta}$ signifies better performance.
@@
We also use the maximum E-measure [8] and the S-measure [7] as validation metrics. E-measure considers both pixel-level errors and image-level errors. S-measure is used to compare the structure similarity of the predicted map and the ground truth with region-aware and object-aware similarity. These two metrics are both better when their values are larger.
@@ 8 Additional ablation experiments
Table 4: The results on ECSSD datasets of ExPert using E-adapter and E-injector with BLIP’s interacted features. L refers to different numbers of cross attention layers. The best performance is in bold.
As shown in Table. 4, one cross attention is enough for the interaction and achieves the best perforance. We assume that the model pre-trained by image-caption pair does not need too much cross attention interaction to highlight the relevant regions in images of text. Besides, less cross attention layers can furthur reduce the size of trained paramters.
@@ 9 Additional qualitative results
Fig. 6 shows some more examples of qualitative results between ExPert and four SOTA models. ExPert can differentiate the salient objects that are similar to or camouflaged in the background, for example, the second row, the third row and the fourth row. Owing to the injection of semantic information, ExPert can handle the ambiguity between different objects in an image, for example, the first row and the fifth row. What’s more, the injection of ViT’s feautre makes up the shallow layers of the full-size scale feature extraction stage of multi-scale backbone. As a result, ExPert can discover some details hard to find. For example the nearly transparent wing of the insect in the sixth row. Fianlly, injection of the complementary features and the semantic information can help ExPert differentiate the shadow distraction in the seventh row where other SOTA models struggles to discern the legs of the man.
Figure 6: The qualitative results of ExPert and four SOTA models. From left to right are the image, the ground truth, ExPert’s masks, EVP’s masks, SelfReformer’s masks, U2Net’s masks and EGNet’s masks. Better visual effect for zooming in.

### v2

@@ 4.1 Experiment Setup
Implemented with PyTorch, ExPert is trained on the DUTS [24] dataset with a batch size of eight using two V100 16G GPUs. The encoder and the decoder are initialized with the publicly released pre-trained weights of SegFormer and the other parameters are initialized randomly. We used the AdamW optimizer and the learning rate is set to 2e-4 with a weight decay of zero.
@@ 4.2.1 Quantitative Comparison
We compare our model’s salient masks with four representative state-of-the art models on five salient object detection datasets: DUTS-TR, DUT-OMRON, ECSSD, HKU-IS and PASCAL-S. Two CNN SOTA models EGNet [34] and U2Net [19] are considered together with two transformer SOTA models SelfReformer [32] and EVP [17]. The metrics are calculated under the same condition using the prediction masks of different models [^6] We use the public codes of SOD_Evaluation_Metrics to compute the metrics.. The prediction masks are all provided by the official release [^7] Considering that EVP’s official mask is 352*352 which is not the original size, we resize the prediction map of EVP to the size of ground truth and then compute the metrics..
The results are shown in Table. 1. Our method achieves the best performance across all the SOTA models on all five datasets, which demonstrates the effectiveness of ExPert and the potential of the transformer backbone on salient object detection. For the MAE metric, ExPert surpasses the second best SelfReformer by 0.0058 (around 21% improvement) in the ECSSD dataset and surpasses EGNet by around 47%. The superiority on all other three metrics demonstrates that ExPert has stronger competence to segment the salient objects in an image. Regarding the size of trained parameters, ExPert is more parameter efficient than all SOTA models except EVP whose trained parameters are also under 100M as ExPert but with a smaller size. Therefore, ExPert realizes a good trade-off between performance and the size of trained parameters. Fig. 4 shows the F-measure curves and the precision-recall curves of ExPert and 4 SOTA models on five datasets. It is observable in these curves that our model consistently outperforms all other models.
Figure 4: The F-measure curves and the precision-recall (PR) curves of ExPert and four SOTA models on five datasets.
To ensure sufficient comparisons, ExPert is also compared to other latest SOD models [^8] Due to the absence of codes or salient maps of some models, we directly use the metrics results in the published paper. The results of M3Net[31] are calculated using the official salient maps of the M3Net SwinB version. including M3Net[31], DSRNet[23], TCRNet[33], BBRF model[18], IMSFNet[26] and CTD-L[12]. As shown in Table. 2, ExPert performs better in three metrics than these latest SOD models in the DUT-OMRON dataset, which demonstrates the superiority of ExPert.
Metrics ExPert M3Net[31] DSRNet[23] TCRNet[33] BBRF[18] IMSFNet[26] CTD-L[12]
[…]
Table 2: The results on the DUT-OMRON dataset of different SOD models and ExPert of metrics MAE, $F_{\beta}$, max E-measure and S-measure. The best results are in bold.
@@ 4.2.2 Qualitative Comparison
Figure 5: The qualitative results of ExPert and four SOTA models. From left to right are the images, the ground truths, ExPert’s masks, EVP’s masks, SelfReformer’s masks, U2Net’s masks and EGNet’s masks. Better visual effect when zooming in.
In Fig. 5, we show the qualitative comparison between our model and SOTA models to give readers an intuitive comprehension. Compared to SOTA models, ExPert’s masks are more accurate in details and can distinguish some ambiguous scenarios like reflection in water in the second row. The head of the girl is similar to the tree in the background in the first row while the hair and the cloth similar to the dark background are confusing in the fifth row. Four SOTA models can’t differentiate these nuances but ExPert can handle the ambiguity. In the first row, ExPert’s mask is even more accurate than the ground truth which contains all the hair. Moreover, the semantic information injection from BLIP aids ExPert in recognizing the relationship between objects in an image. For example, in the fourth row, the baggage on the ground is obviously related to the man. ExPert segments them out while some other SOTA models neglect the baggage or focus wrongly on the street lamp. Another example is the third row, the children in the car should be regarded as a whole with the car. Semantic information also assists ExPert in handling complex scenarios, such as shadow interference or color similarity, for instance, the case of discerning the body of the deer in the shadow. As for salient objects of small sizes, ExPert can well recognize them with clear details and less ambiguity as shown in the last two rows.
@@ 4.3 Ablation study
Tuning methods: To verify the effectiveness of E-adapter on fine-tuning models for salient object detection, the performances of different fine-tuning methods are evaluated on the ECSSD dataset. As shown in Table. 3, full fine-tuning [^9] The full fine-tuning method trains all the parameters of the backbone and decoder using the new datasets. achieves the best performance but requires a large size of trained parameters. Training from scratch requires a large size of trained parameters with poor performances and head tuning [^10] The head tuning method trains only the decoder while keeping the backbone frozen. doesn’t output satisfying results either. As for adapter tuning, the performances are close to the fully tuned version with much fewer trained parameters.
Single-scale and multi-scale backbone: To compare the performance of the single-scale backbone and the multi-scale backbone, all fine-tuning methods are used to fine-tune MiT-B4 SegFormer and two of them are used to fine-tune ViT-B/16. The decoder of ViT is the same as [36], more details can be found in our supplement file. In Table. 3, training from scratch and adapter tuning of ViT are worse than the counterparts of SegFormer because multi-scale features can extract more detailed information in the images. Therefore, ExPert uses multi-scale pre-trained models as the encoder backbone for their finer details.
@@
Adapter level: In [4], the adapter module is side connected inside the transformer block while in ExPert we use the block-level adapter, which is side connected between transformer blocks. In Table. 3, the block-level adapter performs better than the FFT-level adapter which demonstrates that in the segmentation task the block-level adapters are more advantageous. We suppose that the self-attention layers are also important in image recognition and should also be covered by adapters.
Prompt feature: To verify the effectiveness of E-injector and different prompt features, we denote the baseline model as the multi-scale encoder with E-adapter but without E-injector. In Table. 4, we compared E-injector with five prompt features to the baseline. Injecting DINO features as prompt features performs better than the baseline, which shows that object-aware features from DINO can guide the encoder to focus more on salient objects. The ViT features injection is better than the DINO features injection, suggesting that pre-trained models’ semantic information can further boost the performance of SOD. Additionally, although the multi-scale backbone can extract multi-scale salient features, its global layers are usually shallower than those in single-scale backbones. This weakness can be alleviated by the injection of ViT’s features which go through more layers of the full-size scale.
Compared to the injection of ViT’s features, the injection of BLIP’s vision feature performs better. Since BLIP is trained with image-text pairs, the semantic recognition of BLIP is stronger than ViT which is trained with image-label pairs. This indicates that semantic information is critical for the model to detect salient objects. The BLIP+ injection version surpasses the BLIP injection version, verifying that our cross-attention interaction between BLIP’s image features and caption embeddings successfully highlight the salient regions. The best performance comes from the combination of ViT’s features and the BLIP+ features, which captures both the detailed large scale features and the rich semantic information.
@@
Table 4: The results on ECSSD datasets of different prompt features for E-adapter. BLIP+ represents the interacted features after cross-attention of BLIP. The baseline only uses the E-adapter in 3.3 without the E-injector.
@@ 5 Conclusion
Looking ahead, further enhancement of ExPert may include exploration of additional prompt features. It is possible to inject other prompt features such as color or texture information into the backbone. Moreover, ExPert’s paradigm could also be applied to other segmentation tasks, such as semantic segmentation and panoptic segmentation. We also find that ExPert’s performance is influenced by the quality of the generated captions by BLIP. To make ExPert more robust, how to filter out captions of low quality is a challenge. We leave these possible directions for future research.
@@ 6 More details of ExPert
The input image size of our model is 384*384. If the original size is not 384*384, we resize the input images to the certain size. The 384*384 resolution is used for ViT-B/16 [6] and MiT-B4 version of SegFormer [27] for both training and inferring.
We also tried the single-scale pre-trained model ViT-B/16 as the backbone of ExPert. Since the ViT model is trained for classification tasks, the decoder should be changed to a matched one. A simple decoder is adopted for ViT which is drawn from SETR [36]. For $N(=12)$ ViT layers, $N/4$ layers are selected with an interval of 4 as the feature set ${Z_{s}}=\{f_{2},f_{5},f_{8},f_{11}\}$ where $f_{i}$ is the vision feature of the $i_{th}$ layer. Let $\mathcal{U}()$ be the upsampling, $Cat()$ be the concatenate operation applied to all $f_{i}$ in $Z_{s}$ and $\mathcal{S}()$ be the Sigmoid Function. The Mask $M$ is generated as Eq. 8 where $F_{c}$ is the concatenated feature. $Conv2()$ represents two convolution layers with an activation layer and $Conv()$ represents a single convolution layer.
@@ 7 More details of datasets and metrics
Training dataset: We use DUTS [24] as our training dataset following the previous works. DUTS is the largest salient object detection benchmark that contains a training dataset and a test dataset, including respectively 10,553 images and 5,019 images. The training images are all from the ImageNet DET training or validation sets with slight differences in size. We use the test set of DUTS as our validation and evaluation dataset.
Evaluation dataset: We used ECSSD [28], DUT-OMRON [29], HKU-IS [11] and PASCAL-S [14] as the evaluation datasets. The ECSSD dataset contains 1,000 real-world images across different domains with salient objects. DUT-OMRON is a challenging dataset containing 5,168 high quality images with complex backgrounds and various kinds of objects. The HKU-IS dataset has 4,447 well-annotated images characterized by multiple objects in diverse scenarios. PASCAL-S is a subset of the PASCAL VOC segmentation validation dataset, containing 850 images. Unless specified, we used ECSSD as our default evaluation dataset for ablation experiments.
Evaluation metric: Four metrics are adopted for our model evaluation. The first metric is the mean absolute error (MAE) as shown in Eq. 9, which computes the average pixel difference between predicted maps and ground truth labels. A smaller value of MAE signifies better performance.
@@
The second metric is F-measure $F_{\beta}$ [1] which is obtained by leveraging the precision and the recall as shown in Eq. 10. The precision and recall are computed using different thresholds from 0 to 255 between the predicted map and the ground truth. The $\beta^{2}$ is set to 0.3 following the former works. We use the maximum $F_{\beta}$ for the performance of each method. A larger value of $F_{\beta}$ signifies better performance.
@@
We also use the maximum E-measure [8] and the S-measure [7] as validation metrics. E-measure considers both pixel-level errors and image-level errors. S-measure is used to compare the structural similarity of the predicted map and the ground truth with region-aware and object-aware similarity. These two metrics are both better when their values are larger.
@@ 8.1 Cross attention layers
Table 5: The results on the ECSSD datasets of ExPert using E-adapter and E-injector with BLIP’s interacted features. L refers to different numbers of cross attention layers. The best performance is in bold.
As shown in Table. 5, one cross attention is enough for the interaction and achieves the best performance. We assume that the model pre-trained by image-caption pairs does not need too much cross attention interaction to highlight the relevant regions in images of text. Since BLIP is trained with image-caption pairs, the alignment of image and text features is not of pixel scale but rather of patch scale. As auxiliary prompt features, too many cross attentions might lead to unexpected noises that harm the final performance. Besides, fewer cross attention layers can further reduce the size of trained parameters.
8.2 Other transformer backbones
ExPert is backbone-agnostic and can be extended to arbitrary transformer backbones. In the paper, we already showcased that ExPert works well with SegFormer and ViT which represent respectively the multi-scale and single-scale backbone. To further demonstrate ExPert’s flexibility, we replace the SegFormer backbone with the Pyramid Vision Transformer (PVT) backbone[25]. The decoder head shares the same architecture as the multi-scale decoder in Fig. 2 of our paper. As shown in Table. 6, E-adapter and E-injector boost the performance of the PVT backbone compared to full finetuning on SOD datasets with a much smaller size of training parameters which proves the flexibility of ExPert. However, the resolution of the PVT backbone is 224*224 which is smaller than the 384*384 of SegFormer. This might be the reason for the inferior performance of the PVT backbone compared to the SegFormer backbone.
[…]
PVT-F 0.1689 0.7883 0.8396 0.6124 733.73MB
PVT-A 0.0486 0.9213 0.9331 0.8842 50.85MB
PVT-E 0.0409 0.9299 0.9531 0.9026 89.26MB
SegF-A 0.0354 0.9447 0.9626 0.9193 50.9MB
SegF-E 0.0215 0.9550 0.9707 0.9422 80.2MB
Table 6: The results on the ECSSD dataset of the PVT backbone compared to SegFormer backbone. PVT represents the Pyramid Vision Transformer[25] backbone while SegF represents the SegFormer[27] backbone. "-F" represents full finetuning, "-A" represents the adapter tuning with the frozen backbone and only E-adapter without E-injector. "-E" represents the ExPert framework with E-adapter and E-injector. The size column represents the size of trained parameters.
@@ 9 Additional qualitative results
Fig. 6 shows some more examples of qualitative results between ExPert and four SOTA models. ExPert can differentiate the salient objects that are similar to or camouflaged in the background, for example, in the second row, the third row and the fourth row. Owing to the injection of semantic information, ExPert can handle the ambiguity between different objects in an image, for example, in the first row and the fifth row. What’s more, the injection of ViT’s feature makes up for the shallow layers of the full-size scale feature extraction stage of the multi-scale backbone. As a result, ExPert can discover some details hard to find. For example the nearly transparent wing of the insect in the sixth row. Finally, the injection of the complementary features and the semantic information can help ExPert differentiate the shadow distraction in the seventh row where other SOTA models struggle to discern the legs of the man.
Figure 6: The qualitative results of ExPert and four SOTA models. From left to right are the images, the ground truths, ExPert’s masks, EVP’s masks, SelfReformer’s masks, U2Net’s masks and EGNet’s masks. Better visual effect when zooming in.